determine a 95% confidence interval for the mean of a normal population with variance 6.25


1
Expert's answer
2021-03-05T00:45:26-0500

Standart deviation S = "\sqrt{6.25}" = 2.5

From the standard normal (Z-) distribution, z*-value of confidence level 95% is 1.96

"\bar{x} ±1.96\frac{\sigma}{\sqrt{n}}"